OpenStudy (anonymous):

how do you do 2x+16=-x-5

OpenStudy (anonymous):

Take all the x's on one side and the numbers on the other When you take something to the other side the side changes 2x - x = -5-16 x = -21

OpenStudy (anonymous):

add x to both sides then take 16 from both sides then divide both sides by 3

OpenStudy (radar):

I like the second answer better.

OpenStudy (radar):

Which answer is right tiffine?

OpenStudy (anonymous):

Why divide by 3

OpenStudy (anonymous):

im stil confuseddd !

OpenStudy (radar):

Because you have a value for 3x, you want the value for x

OpenStudy (anonymous):

Oh I missed the - in front of the x. So then maths tutor is right

OpenStudy (radar):

Do what mathstutor told you. Everything that was said is correct. Step 1. add x to both sides......what did you get?

OpenStudy (radar):

Then subtract 16 from both sides. What is the result of that?

OpenStudy (radar):

2x+16=-x-5 Now add x to both sides and get: 3x+16=-5

OpenStudy (radar):

Earth to tiffine, are you there?

OpenStudy (radar):

Let me know when you are ready to subtract 16 from both sides.

OpenStudy (radar):

Good luck and study with comprhension.